### v3

# FedAIoT: A Federated Learning Benchmark for Artificial Intelligence of Things

Abstract
There is a significant relevance of federated learning (FL) in the realm of Artificial Intelligence of Things (AIoT). However, most of existing FL works are not conducted on datasets collected from authentic IoT devices that capture unique modalities and inherent challenges of IoT data. To fill this critical gap, in this work, we introduce FedAIoT, an FL benchmark for AIoT. FedAIoT includes eight datasets collected from a wide range of IoT devices. These datasets cover unique IoT modalities and target representative applications of AIoT. FedAIoT also includes a unified end-to-end FL framework for AIoT that simplifies benchmarking the performance of the datasets. Our benchmark results shed light on the opportunities and challenges of FL for AIoT. We hope FedAIoT could serve as an invaluable resource to foster advancements in the important field of FL for AIoT. The repository of FedAIoT is maintained at https://github.com/AIoT-MLSys-Lab/FedAIoT.
Keywords: Benchmarks, Federated Learning, Artificial Intelligence of Things, AIoT
1 Introduction
Internet of Things (IoT) such as smartphones, drones, and sensors deployed at homes are ubiquitous today. The advances in Artificial Intelligence (AI) have boosted the integration of IoT and AI that turns Artificial Intelligence of Things (AIoT) into reality. However, data captured by IoT devices usually contain privacy-sensitive information. In recent years, federated learning (FL) has emerged as a privacy-preserving solution that allows the extraction of knowledge from collected data while keeping the data solely on the devices (Kairouz et al., 2021; Wang et al., 2021; Zhang et al., 2022).
Despite the significant relevance of FL in the AIoT realm, as summarized in Table 9 in Appendix A, most existing FL works are conducted on well-known datasets such as CIFAR-10 and CIFAR-100. These datasets, however, do not originate from authentic IoT devices and thus fail to capture the unique modalities and inherent challenges associated with real-world IoT data. This discrepancy underscores a strong need for an IoT-oriented FL benchmark to fill this critical gap.
In this work, we present FedAIoT, an FL benchmark for AIoT (Figure 1). At its core, FedAIoT includes eight well-chosen datasets collected from a wide range of IoT devices from smartphones, smartwatches, and Wi-Fi routers, to drones, smart home sensors, and head-mounted devices that either have already become an indispensable part of people’s daily lives or are driving emerging applications. These datasets encapsulate a variety of unique IoT-specific data modalities such as wireless data, drone images, and smart home sensor data (e.g., motion, energy, humidity, temperature) that have not been explored in existing FL benchmarks. Moreover, these datasets target some of the most representative applications and innovative use cases of AIoT.
To facilitate benchmarking and ensure reproducibility, FedAIoT includes a unified end-to-end FL framework for AIoT, which covers the complete FL-for-AIoT pipeline: from non-independent and identically distributed (non-IID) data partitioning, IoT-specific data preprocessing, to IoT-friendly models, FL hyperparameters, and IoT-factor emulator. Our framework also includes the implementations of popular schemes, models, and techniques involved in each stage of the FL-for-AIoT pipeline.
Figure 1: Overview of FedAIoT.
We have conducted systematic benchmarking on the eight datasets using the end-to-end framework. Specifically, we examine the impact of varying degrees of non-IID data distributions, FL optimizers, and client sampling ratios on the performance of FL. We also evaluate the impact of noisy labels, a prevalent challenge in IoT datasets, as well as the effects of quantized training, a technique that tackles the practical limitation of resource-constrained IoT devices. Our benchmark results provide valuable information about both the opportunities and challenges of FL for AIoT. Given the significant relevance of FL in the realm of AIoT, we hope FedAIoT could act as a valuable tool to promote advancements in the important area of FL for AIoT. In summary, our work makes the following contributions.
Table 1: Comparison between FedAIoT and existing FL benchmarks.
{tblr}
columneven = c, column3 = c, column5 = c, hline1,10 = -0.08em, hline2 = -, & Data Type FL Framework Designed for IoT Noisy Labels Quantized Training
FLamby Medical Images No No No
FedAudio Audio Data No Uniform  No
FedMultimodal Multimodality Data No Uniform No
FedGraphNN Graph Data No No No
FedNLP Natural Language No No No
FLUTE Images and Text No No Server Side only
FedCV Images No No No
FedAIoT IoT Data Yes Probabilistic Server and Client Side
  • IoT-focused Benchmark. Our benchmark represents the first FL benchmark that focuses on data collected from a diverse set of authentic IoT devices. Moreover, our benchmark includes unique IoT-specific data modalities that previous benchmarks do not include.
  • Unified End-to-End FL Framework for AIoT. We introduce the first unified end-to-end FL framework for AIoT that covers the complete FL-for-AIoT pipeline: from non-IID data partitioning, and IoT-specific data preprocessing, to IoT-friendly models, FL hyperparameters, and IoT-factor emulators.
  • Novel Design of Noisy Labels. Our benchmark also introduces a novel way to design noisy labels. Real-world FL deployments on IoT devices often encounter data labeling errors, which act as noises in federated training. To emulate such noises, different from prior benchmarks that adopt either uniform error distribution or non-uniform error distributions with the assumption that one label can only be mislabeled as another specific label with a random probability, we have designed a new label transition probability matrix based on the insight that labels that are similar to each other are more likely to be mislabeled. This design allows us to simulate data label noises with more realistic transition probabilities and to benchmark FL algorithms for their robustness to such noises.
  • Quantized Training. Lastly, we are also the first FL benchmark to show the effect of quantized training on both server and client sides in the context of FL. Previous benchmarks such as FLUTE (Garcia et al., 2022) demonstrate the effect of quantized training during server-side aggregation for communication reduction. In contrast, our benchmark also incorporates quantized training during client-side training to reduce the memory demands of FL on the device. This is a key difference as IoT devices are often limited by not onlycommunication bandwidth but also on-device memory capacity.
2 Related Work
The importance of data to FL research pushes the development of FL benchmarks on a variety of data modalities. Existing FL benchmarks, however, predominantly center around curating FL datasets in the domain of computer vision (CV) (He et al., 2021b; Garcia et al., 2022), natural language processing (NLP) Lin et al. (2022); Garcia et al. (2022), medical imaging (du Terrail et al., 2022), speech and audio (Zhang et al., 2023; Garcia et al., 2022), and graph neural networks (He et al., 2021a). For example, FedCV He et al. (2021b), FedNLP (Lin et al., 2022), and FedAudio (Zhang et al., 2023) focuses on benchmarking CV, NLP, and audio-related datasets and tasks respectively; FLUTE (Garcia et al., 2022) covers a mix of datasets from CV, NLP, and audio; FLamby (du Terrail et al., 2022) primarily focuses on medical images; and FedMultimodal (Feng et al., 2023) includes multimodal datasets in the domain of emotion recognition, healthcare, multimedia, and social media. As summarized in Table 1, although these benchmarks have contributed to FL research, a dedicated FL benchmark explicitly tailored for IoT data is absent. Compared to these existing FL benchmarks, FedAIoT is specifically designed to fill this critical gap by providing a dedicated FL benchmark on data collected from a wide range of authentic IoT devices.
3 Design of FedAIoT
Table 2: Overview of the datasets included in FedAIoT.
{tblr}
cells = c, column3 = t, column4 = t, hline1-2,10 = -, Dataset & IoT Platform Data Modality Data Dimension Dataset Size # Training Samples # Clients
WISDM-W Smartwatch Accelerometer
Gyroscope $200\times 6$ 294 MB $16,569$ 80
WISDM-P Smartphone Accelerometer
Gyroscope $200\times 6$ 253 MB $13,714$ 80
UT-HAR Wi-Fi Router Wireless Signal $3\times 30\times 250$ 854 MB $3,977$ 20
Widar Wi-Fi Router Wireless Signal $22\times 20\times 20$ 3.3 GB $11,372$ 40
VisDrone Drone Images $3\times 224\times 224$ 1.8 GB $6,471$ 30
CASAS Smart Home Motion Sensor
Door Sensor
Thermostat $2000\times 1$ 233 MB $12,190$ 60
AEP Smart Home Energy, Humidity
Temperature $18\times 1$ 12 MB $15,788$ 80
EPIC-SOUNDS Augmented Reality Acoustics $400\times 128$ 34 GB $60,055$ 210
3.1 Datasets
Objective: The objective of FedAIoT is to provide a benchmark consisting of well-validated and high-quality datasets collected from a wide range of IoT devices, sensor modalities, and applications. We have made significant efforts to examine a much larger pool of existing datasets and select the high-quality ones to include in our benchmark. Table 2 provides an overview of the eight high-quality datasets included in FedAIoT. These datasets have diverse sizes (small: less than 7k samples; medium: 11k to 16k samples; and large: more than 60k samples). The rationale behind this design choice is to accommodate researchers with different computing resources. For example, researchers with limited computing resources can still pick the relatively small datasets in our benchmark to develop and evaluate their algorithms. In this section, we provide a brief overview of each included dataset. More details about the datasets are provided in Appendix B and Table 10.
WISDM: The Wireless Sensor Data Mining (WISDM) dataset (Weiss et al., 2019; Lockhart et al., 2011) is one of the widely used datasets for the task of daily activity recognition using accelerometer and gyroscope sensor data collected from smartphones and smartwatches. WISDM includes data collected from $51$ subjects performing $18$ daily activities, each in a $3$-minute session. We combined activities such as eating soup, chips, pasta, and sandwiches into a single category called “eating”, and removed uncommon activities related to playing with balls, such as kicking, catching, or dribbling. We randomly selected $45$ subjects as the training set and the remaining six subjects were assigned to the test set. Given that the smartwatch data and smartphone data were not collected simultaneously for most subjects and thus were not synchronized precisely, we partition WISDM into two independent datasets: WISDM-W with smartwatch data only and WISDM-P with smartphone data only. The total number of samples in the training and test set is $16,569$ and $4,103$ for WISDM-W and $13,714$ and $4,073$ for WISDM-P respectively. No Licence was explicitly mentioned on the dataset homepage.
UT-HAR: The UT-HAR dataset (Yousefi et al., 2017) is a dataset for contactless activity recognition based on Wi-Fi signals. The Wi-Fi data are in the form of Channel State Information (CSI) collected using three pairs of antennas and an Intel 5300 Network Interface Card (NIC), with each antenna pair capable of capturing $30$ subcarriers of CSI. UT-HAR comprises data collected from subjects performing seven activities such as walking and running. UT-HAR contains a pre-determined training and test set. The total number of training and test samples is $3,977$ and $500$ respectively. No Licence was mentioned on the dataset homepage.
Widar: The Widar dataset (Yang, 2020; Zheng et al., 2019) is designed for contactless gesture recognition using Wi-Fi signal strength measurements collected from strategically placed access points. The data collection system uses an Intel $5300$ NIC with $3\times 3$ antenna pairs. The dataset includes data from $17$ subjects performing $22$ unique gestures like push, pull, sweeping, and clapping. There are some gesture classes that were only contributed by one subject. Including those would risk the model overfitting for that subject instead of the gesture itself. As such, only gestures recorded by more than three subjects are retained, with data from two subjects used during training and the remaining one for the test set. This decision ensures that the test set does not have data from the same subjects as the training set. The resulting dataset includes nine gestures with $11,372$ samples in the training set and $5,222$ in the test set. The dataset is licensed under the Creative Commons Attribution-NonCommercial 4.0 International Licence (CC BY 4).
VisDrone: The VisDrone dataset (Zhu et al., 2021) is a large-scale dataset dedicated to object detection in aerial images captured by drone cameras. VisDrone includes a total of $263$ video clips, which contain $179,264$ frames and $2,512,357$ labeled objects. The labeled objects fall into $12$ categories (e.g., “pedestrian”, “bicycle”, and “car”), recorded under various scenarios such as crowded urban areas, highways, and parks. The dataset contains a pre-determined training and test set. The total number of samples in the training and test set is $6,471$ and $1,610$ respectively. The dataset is licensed under Creative Commons Attribution-NonCommercial-ShareAlike 3.0 License.
CASAS: The CASAS dataset (Schmitter-Edgecombe and Cook, 2009), derived from the CASAS smart home project, is a smart home sensor dataset for the task of recognizing activities of daily living (ADL) based on sequences of sensor states over time to support the application of independent living. Data were collected from three distinct apartments, each equipped with three types of sensors: motion sensors, temperature sensors, and door sensors. We have selected five specific datasets from CASAS named “Milan”, “Cairo”, “Kyoto2”, “Kyoto3”, and “Kyoto4” based on the uniformity of their sensor data representation. The original ADL categories within each dataset have been consolidated into $11$ categories related to home activities such as “sleep”, “eat”, and “bath”. Activities not fitting within these categories were collectively classified as “other”. The training and test set was made using an 80-20 split. Each data sample is a categorical time series of length $2,000$, representing sensor states over a certain period. The total number of samples in the training and test set is $12,190$ and $3,048$ respectively. No Licence was mentioned on the dataset homepage.
AEP: The Appliances Energy Prediction (AEP) dataset (Candanedo et al., 2017) is another smart home sensor dataset but designed for the task of home energy usage prediction. Data were collected from energy sensors, temperature sensors, and humidity sensors installed inside a home every $10$ minutes over $4.5$ months. The number of samples in the training and test set is $15,788$ and $3,947$ respectively. No Licence was mentioned on the dataset homepage.
EPIC-SOUNDS: The EPIC-SOUNDS dataset (Huh et al., 2023) is a large-scale collection of audio recordings for audio-based human activity recognition for augmented reality applications. The audio data were collected from a head-mounted microphone, containing more than $100,000$ categorized segments distributed across $44$ distinct classes. The dataset contains a pre-determined training and test set. The total number of training and test samples is $60,055$ and $40,175$ respectively. The dataset is under CC BY 4 Licence.
Figure 2: Overview of the end-to-end FL framework for AIoT included in FedAIoT.
3.2 End-to-End Federated Learning Framework for AIoT
To benchmark the performance of the datasets and facilitate future research on FL for AIoT, we have designed and developed an end-to-end FL framework for AIoT as another key part of FedAIoT. As illustrated in Figure 2, our framework covers the complete FL-for-AIoT pipeline, including (1) non-IID data partitioning, (2) IoT-specific data preprocessing, (3) IoT-friendly models, (4) FL hyperparameters, and (5) IoT-factor emulator. In this section, we describe these components in detail.
3.2.1 Non-IID Data Partitioning
A key characteristic of FL is that data distribution at different clients is non-IID. The objective of non-IID data partitioning is to partition the training set such that data allocated to different clients follow the non-IID distribution. The eight datasets included in FedAIoT cover three fundamental tasks: classification, regression, and object detection. As summarized in Table 3, FedAIoT incorporates three different non-IID data partitioning schemes designed for the three tasks respectively.
Table 3: Non-IID data partitioning schemes and models used for each dataset.
{tblr}
cells = c, cell21 = r=2, hline1,5 = -0.08em, hline2 = -, Dataset & WISDM-W WISDM-P UT-HAR Widar VisDrone CASAS AEP EPIC-SOUNDS
Partition
Output Output Output Output Input Output Output Output
Labels Labels Labels Labels Features Labels Distribution Labels
Model LSTM LSTM ResNet18 ResNet18 YOLOv8n BiLSTM MLP ResNet18
Scheme#1: Non-IID Partition over Output Labels. For the task of classification (WISDM-W, WISDM-P, UT-HAR, Widar, CASAS, EPIC-SOUNDS) with $C$ classes, we first generate a distribution over the classes for each client by drawing from a Dirichlet distribution (formal definition in Appendix C) with parameter $\alpha$ (Hsu et al., 2019), where lower values of $\alpha$ generate more skewed distribution whereas higher values of $\alpha$ result in more balanced class distributions. We use the same $\alpha$ to determine the number of samples each client contains. In addition, by drawing from a Dirichlet distribution with parameter $\alpha$, we create a distribution over the total number of samples, which is then used to allocate a varying number of samples to each client, where lower values of $\alpha$ lead to a few clients holding a majority of the samples whereas higher values of $\alpha$ create a more balanced distribution of samples across clients. Therefore, this approach allows us to generate non-IID data partitions such that both class distribution and the number of samples can vary across the clients.
Scheme#2: Non-IID Partition over Input Features. The task of object detection (VisDrone) does not have specific classes. In such case, we use the input features to create non-IID partitions. Specifically, similar to He et al. (2021b), we first used ImageNet (Russakovsky et al., 2015) features generated from the VGG19 model (Liu and Deng, 2015), which encapsulate visual information required for subsequent analysis. With these ImageNet features as inputs, we performed clustering in the feature space using $K$-nearest neighbors to partition the dataset into clusters. Each cluster is a pseudo-class, representing images sharing common visual characteristics as per the extracted ImageNet features. Lastly, Dirichlet allocation was applied on top of the pseudo-classes to create the non-IID distribution across different clients.
Scheme#3: Non-IID Partition over Output Distribution. For the regression task (AEP) where output is characterized as a continuous variable, we utilize Quantile Binning (Pyle, 1999). Specifically, we divide the range of the output variable into equal groups or quantiles, ensuring that each bin accommodates roughly the same number of samples. Each category or bin is treated as a pseudo-class. Note that the number of quantiles can be set to any value in our framework. We used ten as an example to demonstrate the results. Lastly, we apply Dirichlet allocation to generate the non-IID distribution of data across the clients.
3.2.2 IoT-specific Data Preprocessing
The eight datasets included in FedAIoT cover diverse IoT data modalities such as wireless signals, drone images, and smart home sensor data. FedAIoT incorporates a suite of IoT-specific data preprocessing techniques that are designed for such IoT data modalities accordingly.
WISDM: We followed the standard preprocessing techniques used in accelerometer and gyroscope-based activity recognition for WISDM (Ravi et al., 2005; Reyes-Ortiz et al., 2016; Ronao and Cho, 2016). Specifically, for each $3$-minute session, we used a $10$-second sliding window with $50$% overlap to extract samples from the raw accelerometer and gyroscope data sequences. We then normalize each dimension of the extracted samples by removing the mean and scaling to unit variance.
UT-HAR: We followed Yang et al. (2023) and applied a sliding window of $250$ packets with $50$% overlap to extract samples from the raw Wi-Fi data from all three antennas. We then normalize each dimension of the extracted samples by removing the mean and scaling to unit variance.
Widar: We first adopted the body velocity profile (BVP) processing technique as outlined in Yang et al. (2023); Zheng et al. (2019) to handle environmental variations from the data. We then applied standard scalar normalization to normalize the data. This creates data samples with the shape of $22\times 20\times 20$ reflecting time axis, $x$, and $y$ velocity features respectively.
VisDrone: We first normalized the pixel values of drone images to range from 0 to 1. Data augmentation techniques including random shifts in Hue, Saturation, and Value color space, image compression, shearing transformations, scaling transformations, horizontal and vertical flipping, and MixUp were applied to increase the diversity of the dataset.
CASAS: We followed Liciotti et al. (2019) to transform the sensor readings into categorical sequences, creating a form of semantic encoding. Every possible temperature setting is assigned a distinct categorical value. Individual instances of motion and door sensor activation (on or off) are also assigned a categorical value. Subsequently, for a sensor activation, we extract a sequence consisting of the previous $2,000$ sensor activations which we then use for activity prediction.
[…]
EPIC-SOUNDS: We first performed a Short-Time Fourier Transform (STFT) on raw audio data followed by applying a Hanning window of 10ms duration and a step size of 5ms to ensure appropriate spectral resolution. We then extracted $128$ Mel spectrogram features. To further enhance the data, we applied a natural logarithm scaling to the Mel spectrogram output. Lastly, we padded each segment to reach a consistent length of $400$.
3.2.3 IoT-friendly Models
Given that FedAIoT focuses on resource-constrained IoT devices, our choice of models is informed by a combination of model accuracies and model efficiency (Reinbothe, 2023; Yang et al., 2023; Terven et al., 2023; Liciotti et al., 2019; Seyedzadeh et al., 2018; Sholahudin et al., 2016; Huh et al., 2023). For each included dataset, we evaluated multiple model candidates, and selected the best-performing ones that adhere to the resource constraints of representative IoT devices listed in Table 4. As an example, for UT-HAR dataset, two model candidates (ViT and ResNet18) have similar accuracy, and we selected ResNet18 in our benchmark as it has better model efficiency. Table 3 lists the selected model for each dataset. The detail of the architecture of each model is described in Appendix D.
Table 4: Representative IoT devices for each dataset.
{tblr}
cells = c, cell22 = r=3, hline1-2,5-10 = -, Dataset & Application IoT Platform Representative Devices Hardware RAM Size
WISDM-W Activity Recognition Smartwatch Apple Watch 8 512 MB to 1 GB
WISDM-P Smartphone iPhone 14 6 GB
UT-HAR Wi-Fi Router TP-Link AX1800 64 MB to 1 GB
Widar Gesture Recognition Wi-Fi Router TP-Link AX1800 64 MB to 1 GB
CASAS Independent Living Smart Home Raspberry Pi 4 1 GB to 8 GB
AEP Energy Prediction Smart Home Raspberry Pi 4 1 GB to 8 GB
VisDrone Objective Detection Drone Dji Mavic 3 + Raspberry Pi 4 1 GB to 8 GB
EPIC-SOUNDS Augmented Reality Head-mounted Device GoPro / AR Headset 1 GB to 8 GB
3.2.4 FL Hyperparameters
Data Heterogeneity Level. Data heterogeneity (i.e., non-IIDness) is a fundamental challenge in FL. As outlined in Section 3.2.1, FedAIoT facilitates the creation of diverse non-IID data partitions, which enables the simulation of different data heterogeneity levels to meet experiment requirements.
FL Optimizer. FedAIoT supports a handful of commonly used FL optimizers. In the experiment section, we showcase the benchmark results of two of the most commonly used FL optimizers: FedAvg (McMahan et al., 2017) and FedOPT (Reddi et al., 2021).
Client Sampling Ratio. Client sampling ratio denotes the proportion of clients selected for local training in each round of FL. This hyperparameter plays a crucial role as it directly influences the computation and communication costs associated with FL. FedAIoT facilitates the creation of diverse client sampling ratios and the evaluation of its impact on both model performance and convergence speed during federated training.
Details on other hyperparameters used for each experiment are described in Appendix E.
3.2.5 IoT-Factor Emulator
Noisy Labels. In real-world scenarios, IoT sensor data can be mistakenly labeled. These labels introduce noise into the federated training process. Therefore, FL systems need to be designed with robustness to label noise as a crucial feature. Label noise is a well-discussed concept in centralized training with established methods to mitigate it (Song et al., 2023). In general, label noise can be emulated in two ways: Symmetric Noise where the injected noise is uniformly distributed across the labels (Tanaka et al., 2018); and Asymmetric Noise injects noise non-uniformly across different labels using a label transition probability matrix (Patrini et al., 2017; Tanaka et al., 2018). In the context of FL, Xu et al. (2022); Zhang et al. (2023); Feng et al. (2023) propose to use symmetric noise to simulate label noise, evaluating FL algorithms at various noise levels. Fang and Ye (2022); Kim et al. (2022); Wu et al. (2023); Liang et al. (2024); Tsouvalas et al. (2024), on the other hand, propose to use asymmetric noise to simulate label noise. In particular, Fang and Ye (2022); Kim et al. (2022); Liang et al. (2024); Tsouvalas et al. (2024) design label transition probabilities in a pairwise manner, assuming that one label can only be mislabeled as another specific label with a random probability. In contrast, in FedAIoT, we propose a new label transition probability matrix that breaks this assumption, where a label can be mistakenly labeled as any other label with a learned probability. Specifically, we augment the ground truth labels of a dataset with a label transition probability matrix $Q$ where, $Q_{ij}$ is the probability that the true label $i$ is changed to a different label $j$, i.e., $P(\hat{y}=j\mid y=i)$. The label transition probability matrix was constructed based on centralized training results, where the elements of $Q_{ij}$ was determined from the ratio of samples labeled as $j$ by a centrally trained model to those with ground truth label $i$.
Quantized Training. IoT devices are resource-constrained. Previous benchmarks such as FLUTE (Garcia et al., 2022) examined the performance of quantized training during server-side model aggregation to reduce the communication cost of FL. Ozkara et al. (2021) and Gupta et al. (2023) study quantization-aware training but still use full precision during training. In contrast, FedAIoT incorporates not only quantized model aggregation at the server side but also quantized training at the client side to reduce the memory demands of FL on client IoT devices. Quantized training on both server and client sides is key to enabling FL for AIoT as IoT devices are not just restrained in communication bandwidth but also in on-device memory.
4 Benchmark Results and Analysis
We implemented FedAIoT using PyTorch (Paszke et al., 2019) and Ray (Moritz et al., 2018) and conducted our experiments on NVIDIA A6000 GPUs. We run each of our experiments using three random seeds and report the mean and standard deviation.
4.1 Overall Performance
First, we benchmark the FL performance under two FL optimizers, FedAvg and FedOPT, under low ($\alpha=0.5$) and high ($\alpha=0.1$) data heterogeneity levels, and compare it against centralized training.
Table 5: Overall performance.
Dataset Metric Centralized Low Data Heterogeneity ($\mathbf{\alpha=0.5}$) High Data Heterogeneity ($\mathbf{\alpha=0.1}$)
FedAvg FedOPT FedAvg FedOPT
WISDM-W Accuracy (%) $74.05\pm 2.47$ $70.03\pm 0.13$ $71.50\pm 1.52$ $68.51\pm 2.21$ $65.76\pm 2.42$
WISDM-P Accuracy (%) $36.88\pm 1.08$ $36.21\pm 0.19$ $34.32\pm 0.84$ $34.28\pm 3.28$ $32.99\pm 0.55$
UT-HAR Accuracy (%) $95.24\pm 0.75$ $94.03\pm 0.63$ $94.10\pm 0.84$ $74.24\pm 3.87$ $87.78\pm 5.48$
Widar Accuracy (%) $61.24\pm 0.56$ $59.21\pm 1.79$ $56.26\pm 3.11$ $54.76\pm 0.42$ $47.99\pm 3.99$
VisDrone MAP-50 (%) $34.26\pm 1.56$ $32.70\pm 1.19$ $32.21\pm 0.28$ $31.23\pm 0.70$ $31.51\pm 2.18$
CASAS Accuracy (%) $83.70\pm 2.21$ $75.93\pm 2.82$ $76.40\pm 2.20$ $74.72\pm 1.32$ $75.36\pm 2.40$
AEP $R^{2}$ $0.586\pm 0.006$ $0.502\pm 0.024$ $0.503\pm 0.011$ $0.407\pm 0.003$ $0.475\pm 0.016$
EPIC-SOUNDS Accuracy (%) $46.97\pm 0.24$ $45.51\pm 1.07$ $42.39\pm 2.01$ $33.02\pm 5.62$ $37.21\pm 2.68$
Benchmark Results: Table 5 summarizes our results. We make three observations. (1) Data heterogeneity level and FL optimizer have different impacts on different datasets. In particular, the performance of UT-HAR, AEP, and EPIC-SOUNDS is extremely sensitive to the data heterogeneity level. In contrast, WISDM-P, CASAS, and VisDrone show limited accuracy differences under different data heterogeneity levels. (2) Under high data heterogeneity, FedAvg has a better performance compared to FedOPT in WISDM and Widar datasets and has lower deviation for all datasets except WISDM-P and EPIC-SOUNDS. The performance gap between the two FL optimizers reduces under low data heterogeneity. (3) Compared to the other datasets, CASAS, AEP, and WISDM-W have higher accuracy margins between centralized training and low data heterogeneity.
4.2 Impact of Client Sampling Ratio
Second, we benchmark the FL performance under two client sampling ratios: $10$% and $30$%. We report the maximum accuracy reached after completing $50$%, $80$%, and $100$% of the total training rounds for both these ratios under high data heterogeneity, thereby offering empirical evidence of how the model performance and convergence rate are impacted by the client sampling ratio.
Benchmark Results: Table 6 summarizes our results. We make two observations. (1) When the client sampling ratio increases from $10$% to $30$%, the model performance increases at $50$%, $80$%, and $100$% of the total training rounds across all the datasets. This demonstrates the importance of the client sampling ratio to the model performance. (2) However, a higher sampling ratio does not always speed up model convergence to the same extent. For example, model convergence of CASAS and VisDrone are comparable at both sampling ratios whereas it is much faster for UT-HAR and AEP.
4.3 Impact of Noisy Labels
Next, we examine the impact of noisy labels on the FL performance under two label error ratios: $10$% and $30$%, and compare these results with the control scenario that involves no label errors. Note that we only showcase this for WISDM, UT-HAR, Widar, CASAS, and EPIC-SOUNDS as these are classification tasks, and the concept of noisy labels only applies to classification tasks.
Table 6: Impact of client sampling ratio.
Dataset Training Rounds Low Client Sampling Ratio ($10\%$) High Client Sampling Ratio ($30\%$)
50% Rounds 80% Rounds 100% Rounds 50% Rounds 80% Rounds 100% Rounds
WISDM-W $400$ $58.81\pm 1.43$ $63.82\pm 1.53$ $68.51\pm 2.21$ $65.57\pm 2.10$ $67.23\pm 0.77$ $69.21\pm 1.13$
WISDM-P $400$ $29.49\pm 3.65$ $31.65\pm 1.42$ $34.28\pm 3.28$ $33.73\pm 2.77$ $34.01\pm 2.27$ $36.01\pm 2.23$
UT-HAR $2000$ $61.81\pm 7.01$ $70.76\pm 2.23$ $74.24\pm 3.87$ $86.46\pm 10.90$ $90.84\pm 4.42$ $92.51\pm 2.65$
Widar $1500$ $47.55\pm 1.20$ $50.65\pm 0.24$ $54.76\pm 0.42$ $53.93\pm 2.90$ $55.74\pm 2.15$ $57.39\pm 3.14$
VisDrone $600$ $27.07\pm 3.09$ $31.05\pm 1.55$ $31.23\pm 0.70$ $30.56\pm 2.71$ $33.52\pm 2.90$ $34.85\pm 0.83$
CASAS $400$ $71.68\pm 1.96$ $74.19\pm 1.26$ $74.72\pm 1.32$ $73.89\pm 1.16$ $74.68\pm 1.50$ $76.12\pm 2.03$
AEP $3000$ $0.325\pm 0.013$ $0.371\pm 0.017$ $0.407\pm 0.003$ $0.502\pm 0.006$ $0.523\pm 0.014$ $0.538\pm 0.005$
EPIC-SOUNDS $300$ $20.99\pm 5.19$ $25.73\pm 1.99$ $28.89\pm 2.82$ $23.70\pm 6.25$ $31.74\pm 7.83$ $35.11\pm 1.99$
Table 7: Impact of noisy labels.
{tblr}
cells = c, hline1,5 = -0.08em Noisy Label Ratio & WISDM-W WISDM-P UT-HAR Widar CASAS EPIC-SOUNDS
0% $68.51\pm 2.21$ $34.28\pm 3.28$ $74.24\pm 3.87$ $54.76\pm 0.42$ $74.72\pm 1.32$ $28.89\pm 2.82$
10% $50.63\pm 4.19$ $28.85\pm 1.44$ $73.75\pm 5.67$ $34.03\pm 0.33$ $65.01\pm 2.98$ $21.43\pm 3.86$
30% $47.90\pm 3.05$ $27.68\pm 0.39$ $70.55\pm 3.27$ $27.20\pm 0.56$ $63.16\pm 1.34$ $13.30\pm 0.42$
Benchmark Results: Table 7 summarizes our results. We make two observations. (1) As the ratio of erroneous labels increases, the performance of the models decreases across all the datasets, and the impact of noisy labels varies across different datasets. For example, UT-HAR only experiences a little performance drop at $10$% label error ratio, but its performance drops more when the label error ratio increases to $30$%. (2) In contrast, WISDM, Widar, CASAS, and EPIC Sounds are very sensitive to label noise and show significant accuracy drop even at $10$% label error ratio.
4.4 Performance on Quantized Training
Lastly, we examine the impact of quantized training on FL under half-precision (FP16) [^1] PyTorch does not support lower quantization levels like INT8 and INT4 during training as referenced Pytorch (2023) at the time of writing and hence those were excluded. Additionally, FP8 is only supported for H100 GPUs, not the mobile GPUs used for IoT devices.. We assess model accuracy and memory usage under FP16 and compare the results to those from the full-precision (FP32) models. Memory usage is measured by analyzing the GPU memory usage of a model when trained with the same batch size under a centralized setting. Note that we use memory usage as the metric since it is a relatively consistent and hardware-independent metric. In contrast, other metrics such as computation speed and energy are highly hardware-dependent. Depending on the chipset that the IoT devices use, the computation speed and energy could exhibit wide variations. More importantly, new and more advanced chipsets are produced every year. The updates of the chipsets would inevitably make the benchmarking results quickly obsolete and out of date.
Benchmark Results: Table 8 summarizes the model performance and memory usage at two precision levels. We make three observations: (1) As expected, the memory usage significantly decreases when using FP16 precision, ranging from $57.0$% to $63.3$% reduction across different datasets. (2) Similar to Micikevicius et al. (2018), we also observe that model performance associated with the precision levels varies depending on the dataset. For AEP and EPIC-SOUNDS, the FP16 models improve the performance compared to the FP32 models. (3) Widar and WISDM-W have a significant decline in performance when quantized to FP16 precision.
Table 8: Performance on quantized training.
Dataset Metric FP32 FP16
Model Performance Memory Usage Model Performance Memory Usage
WISDM-W Accuracy (%) $68.51\pm 2.21$ 1444 MB $60.31\pm 5.38$ 564 MB ($\downarrow$ 60.9%)
WISDM-P Accuracy (%) $34.28\pm 3.28$ 1444 MB $30.22\pm 2.05$ 564 MB ($\downarrow$ 60.9%)
UT-HAR Accuracy (%) $74.24\pm 3.87$ 1716 MB $72.86\pm 4.49$ 639 MB ($\downarrow$ 62.8%)
Widar Accuracy (%) $54.76\pm 0.42$ 1734 MB $34.03\pm 0.33$ 636 MB ($\downarrow$ 63.3%)
VisDrone MAP-50 (%) $31.23\pm 0.70$ 8369 MB $29.17\pm 4.70$ 3515 MB ($\downarrow$ 60.0%)
CASAS Accuracy (%) $74.72\pm 1.32$ 1834 MB $72.86\pm 4.49$ 732 MB ($\downarrow$ 60.1%)
AEP $R^{2}$ $0.407\pm 0.003$ 1201 MB $0.469\pm 0.044$ 500 MB ($\downarrow$ 58.4%)
EPIC-SOUNDS Accuracy (%) $33.02\pm 5.62$ 2176 MB $35.43\pm 6.61$ 936 MB ($\downarrow$ 57.0%)
4.5 Insights from Benchmark Results
Need for Resilience on High Data Heterogeneity: As shown in Table 5, datasets can be sensitive to data heterogeneity. We observe that UT-HAR, Widar, AEP, and EPIC-SOUNDS show a significant impact under high data heterogeneity. These findings emphasize the need for developing advanced FL algorithms for data modalities that are sensitive to high data heterogeneity. Handling high data heterogeneity is still an open question in FL (Zhao et al., 2018; Li et al., 2020a) and our benchmark shows that it is a limiting factor for some of the IoT data modalities as well. To mitigate high data heterogeneity, although many techniques have been proposed (Sattler et al., 2019; Arivazhagan et al., 2019), their performance is still constrained. Most recently, some exploration on incorporating generative pre-trained transformers (GPT) as part of the FL framework has shown great performance in mitigating high data heterogeneity (Zhang et al., 2024).
Need for Balancing between Client Sampling Ratio and Resource Consumption of IoT Devices: Table 6 reveals that a higher sampling ratio can lead to improved performance in the long run. However, higher client sampling ratios generally entail increased communication overheads and energy consumption, which may not be desirable for IoT devices. Therefore, it is crucial to identify a balance between the client sampling ratio and resource consumption.
Need for Resilience on Noisy Labels: As shown in Table 7, certain datasets exhibit high sensitivity to label errors, significantly deterring FL performance. Notably, both WISDM-W and Widar experience a drastic decrease in accuracy when faced with a $10$% label noise ratio. Given the inevitability of noise in real FL deployments where private data is unmonitored except by the respective data owners, the development of label noise resilient techniques becomes crucial for achieving reliable FL performance. However, handling noisy labels is a much less explored topic in FL. Pioneering work on this topic is still exploring the effects of noisy labels on the performance of FL (Zhang et al., 2023; Feng et al., 2023). To mitigate the effect of noisy labels, for low label noise rates, techniques such as data augmentation and regularization have been shown to be effective in centralized training settings (Shorten and Khoshgoftaar, 2019; Zhang et al., 2018). If the label error rate is high, techniques such as knowledge distillation (Hinton et al., 2015), mixup (Zhang et al., 2018), and bootstrapping (Reed et al., 2015) have been proposed. We expect such techniques or their variants would help in the FL setting.
Need for Quantized Training: Table 4 highlights the need for quantized training given the limited RAM resources on representative IoT devices. From our results summarized in Table 8, we observe that FP16 quantization does not affect accuracy drastically unless normalization layers are present, like batch normalization (Jacob et al., 2018). To mitigate this issue, techniques such as incremental quantization (Zhou et al., 2017) or ternary weight networks (Liu et al., 2023) have shown great performance in the centralized training setting. We expect such techniques or their variants would help in the FL setting as well.
5 Conclusion
In this paper, we presented FedAIoT, an FL benchmark for AIoT. FedAIoT includes eight datasets collected from a wide range of authentic IoT devices as well as a unified end-to-end FL framework for AIoT that covers the complete FL-for-AIoT pipeline. We have benchmarked the performance of the datasets and provided insights on the opportunities and challenges of FL for AIoT. Moving forward, we aim to foster community collaboration by launching an open-source repository for this benchmark, continually enriching it with more datasets, algorithms, and thorough analytical evaluations, ensuring it remains a dynamic and invaluable resource for FL for AIoT research.
6 Acknowledgement
We would like to thank the action editor Peter Mattson and the anonymous reviewers of the Journal of Data-centric Machine Learning Research for their helpful and constructive comments. Zhichao Cao is supported in part by National Science Foundation (NSF) under award NeTS-2312675. JeongGil Ko is partially supported by the Ministry of Science and ICT (MSIT) and the IITP (IITP-2022-2020-001461, IITP-2022-0-00420) in the Republic of Korea. Tiantian Feng and Shrikanth S. Narayanan are supported in part by USC + Amazon Center on Secure & Trusted ML. Tuo Zhang and Salman Avestimehr are supported in part by ARO award W911NF1810400, ONR Award No. N00014-16-1-2189 and NSF under NSF SATC ProperData Center. Samiul Alam and Mi Zhang are supported in part by the Meta Reality Labs Faculty Research Award. The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.
Appendix A Datasets used in Existing Representative FL Works
Table 9: Summary of the datasets used in existing representative FL works.
{tblr}
width = colspec = Q[450]Q[781], column2 = l, hline1,15 = -0.08em, hline2 = -,                  Method &                          Datasets Used
FedAvg (McMahan et al., 2017) MNIST (LeCun et al., 1998), Shakespeare (Karpathy, 2015)
FedOPT (Reddi et al., 2021) CIFAR-10/100 (Krizhevsky et al., 2009), EMNIST (Cohen et al., 2017), Shakespeare (Karpathy, 2015), StackOverflow (TensorFlow Federated, 2019)
Scaffold (Karimireddy et al., 2020) EMNIST (Cohen et al., 2017)
FedProx (Li et al., 2020b) MNIST(LeCun et al., 1998), FEMNIST (Xiao et al., 2017), Shakespeare (Karpathy, 2015), Sent 140 (Goel et al., 2016)
FedDF (Lin et al., 2020) CIFAR-10/100 (Krizhevsky et al., 2009), AG News (Zhang et al., 2015), SST2 (Socher et al., 2013)
DS-FL (Itahara et al., 2021) MNIST (LeCun et al., 1998), FEMNIST (Xiao et al., 2017)
Fed-ET(Cho et al., 2022) CIFAR-10/100 (Krizhevsky et al., 2009), Sent 140 (Goel et al., 2016)
HeteroFL (Diao et al., 2021) MINIST (LeCun et al., 1998), CIFAR-10/100 (Krizhevsky et al., 2009), WikiText (Merity et al., 2017)
FedRolex (Alam et al., 2022) CIFAR-10/100 (Krizhevsky et al., 2009), StackOverflow (TensorFlow Federated, 2019)
FedGen (Venkateswaran et al., 2023) SST-2 (Socher et al., 2013), AG News (Zhang et al., 2015), UT-HAR (Yousefi et al., 2017), Water Quality (Environmental Protection Agency, 2021), BPIC 2018 (van Dongen and Borchert, 2018)
DynaFed (Pi et al., 2023) FMNIST (Xiao et al., 2017), CIFAR-10 (Krizhevsky et al., 2009), CINIC-10 (Darlow et al., 2018)
GPT-FL (Zhang et al., 2024) CIFAR-10/100 (Krizhevsky et al., 2009) , Flower-102 (Lv et al., 2021), Google Command (Warden, 2018), ESC-50 (Piczak, 2015)
pFedHR (Wang et al., 2024) MNIST (LeCun et al., 1998), SVHN (Yang et al., 2021), and CIFAR-10 (Krizhevsky et al., 2009)
Appendix B Dataset Details
B.1 WISDM
The WISDM dataset comprises raw accelerometer and gyroscope data collected from 51 subjects performing 18 activities for three minutes each. Data were gathered at a 20Hz sampling rate from both a smartphone (Google Nexus 5/5x or Samsung Galaxy S5) and a smartwatch (LG G Watch). Data for each device and sensor type are stored in different directories, resulting in four directories overall. Each directory contains 51 files, each corresponding to a subject. The data entry format is: <subject-id, activity-code, timestamp, x, y, z>. Separate files for the gyroscope and accelerometer readings are provided and are later combined by matching timestamps. Subject ID is given from 1600 to 1650 and the activity code is an alphabetical character between ‘A’ and ‘S’ excluding ‘N’. The timestamp is in Unix time. The code to read and partition the data into 10s segments is provided by our benchmark. The input shape of the processed data is $200\times 6$. The original dataset is available at https://archive.ics.uci.edu/dataset/507/wisdm+smartphone+and+smartwatch+activity+and+biometrics+dataset.
B.2 UT-HAR
The UT-HAR dataset was collected using the Linux 802.11n Channel State Information (CSI) Tool for the task of Human Activity Recognition (HAR). The original data consist of two file types: “input" and “annotation". “input" files contain Wi-Fi CSI data. The first column indicates the timestamp in Unix. Columns 2-91 represent amplitude data for 30 subcarriers across three antennas, and columns 92-181 contain the corresponding phase information. “annotation" files provide the corresponding activity labels, serving as the ground truth for HAR. In our benchmark, only amplitude is used. The final samples are created by taking a sliding window of size $250$ where each sample consists of amplitude information across three antennas and from 30 subcarriers and has shape $3\times 30\times 250$. The original dataset is available at https://github.com/ermongroup/Wifi_Activity_Recognition/tree/master.
B.3 Widar
The Widar dataset (Widar3.0) was collected with a system comprising one transmitter and three receivers, all equipped with Intel 5300 wireless NICs. The system uses the Linux CSI Tool to record the Wi-Fi data. Devices operate in monitor mode on channel 165 at 5.825 GHz. The transmitter broadcasts $1,000$ Wi-Fi packets per second while receivers capture data using their three linearly arranged antennas. In our benchmark, we use the processed body velocity profile (BVP) features extracted from the dataset. The size of each data sample after processing is $22\times 20\times 20$ consisting of $22$ samples over time each having $20$ BVP features each in both $x$ and $y$ directions. The raw dataset is available for download at http://tns.thss.tsinghua.edu.cn/widar3.0/index.html.
B.4 VisDrone
The VisDrone dataset was collected by the AISKYEYE team at Tianjin University, China. It comprises 288 video clips with 261,908 frames and 10,209 static images captured by cameras mounted on drones at 14 different cities in China in diverse environments, scenarios, weather, and lighting conditions. The frames were manually annotated with over 2.6 million bounding boxes of common targets like pedestrians, cars, and bicycles. Additional attributes like scene visibility, object class, and occlusion are also provided for enhanced data utilization. The dataset is available at https://github.com/VisDrone/VisDrone-Dataset.
B.5 CASAS
The CASAS dataset is a collection of data generated in smart home environments, where intelligent software uses sensors deployed at homes to monitor resident activities and conditions within the space. The CASAS project considers environments as intelligent agents and employs custom IoT hardware known as Smart Home in a Box (SHiB), which encompasses the necessary sensors, devices, and software. The sensors in SHiB perceive the status of residents and their surroundings, and through controllers, the system acts to enhance living conditions by optimizing comfort, safety, and productivity. The CASAS dataset includes the date (in yyyy-mm-dd format), time (in hh:mm:ss.ms format), sensor name, sensor readings, and an activity label in string format. The data were collected in real-time as residents go about their daily activities. The code to extract categorical sensor readings to create input sequences and labels is provided in our benchmark. The CASAS dataset can be downloaded from https://casas.wsu.edu/datasets/.
B.6 AEP
The AEP dataset, collected over 4.5 months, comprises readings taken every 10 minutes from a ZigBee wireless sensor network monitoring house temperature and humidity. Each wireless node transmitted data around every 3.3 minutes, which were then averaged over 10-minute periods. Additionally, energy data was logged every 10 minutes via m-bus energy meters. The dataset includes attributes such as date and time (in year-month-day hour:minute:second format), the energy usage of appliances and lights (in Wh), temperature and humidity in various rooms including the kitchen ($T_{1}$, $RH_{1}$), living room ($T_{2}$, $RH_{2}$), laundry room ($T_{3}$, $RH_{3}$), office room ($T4$, $RH_{4}$), bathroom ($T_{5}$, $RH_{5}$), ironing room ($T_{7}$, $RH_{7}$), teenager room ($T_{8}$, $RH_{8}$), and parents room ($T_{9}$, $RH_{9}$), and temperature and humidity outside the building ($T_{6}$, $RH_{6}$) - all with temperatures in Celsius and humidity in percentages. Additionally, weather data from Chievres Airport, Belgium was incorporated, consisting of outside temperature (To in Celsius), pressure (in mm Hg), humidity ($RH_{out}$ in %), wind speed (in m/s), visibility (in km), and dew point ($T_{dewpoint}$ in °C). The dataset is available at https://archive.ics.uci.edu/dataset/374/appliances+energy+prediction.
B.7 EPIC-SOUNDS
As an extension of the EPIC-KITCHENS-100 dataset, the EPIC-SOUNDS dataset focuses on annotating distinct audio events in the videos of EPIC-KITCHENS-100. The annotations include the time intervals during which each audio event occurs, along with a text description explaining the nature of the sound. Given the variation in video lengths in the dataset, which range from 30 seconds to 1.5 hours, the videos are segmented into clips of 3-4 minutes each to make the annotation process more manageable. In order to ensure that annotators concentrate solely on the audio aspects, only the audio stream is provided to them. This decision is taken to prevent bias that could be introduced by the visual and contextual elements in the videos. Additionally, annotators are given access to the plotted audio waveforms. These visual representations of the audio data help the annotators by guiding them in pinpointing specific sound patterns, thus making the annotation process more efficient and targeted. The EPIC-SOUNDS dataset can be extracted from the EPIC-KITHENS-100 dataset with the GitHub repo at https://github.com/epic-kitchens/epic-sounds-annotations. The extracted audio data in the form of HDF5 file format can also be requested from mailto:uob-epic-kitchens@bristol.ac.uk.
Table 10: Details of the datasets included in FedAIoT.
{tblr}
width = colspec = Q[90]Q[325]Q[131]Q[130]Q[279], column1-5 = c, hline1,9 = -0.08em, hline2 = -, Dataset & Source Device #Subjects Involved Conditions Labeling
WISDM Smartphone (Google Nexus 5/5x or Samsung Galaxy S5)
Smartwatch (LG G Watch) 51 University Lab Each sample labeled using app by observer
UT-HAR Intel 5300 NIC Receiver 10 Indoor Office Labeled by comparing with video feed by observer
WIDAR Intel 5300 NIC Receiver 17 University Lab Labeled manually by external observer
VisDrone DJI Mavic Phantom Series Drone N/A 14 Cities in China Crowdsourced labeling
CASAS Smart Home in a Box Suite 20 Apartment Labeled by participants
AEP Zigbee wireless sensor network
Energy Meter 4 2 Storey House Target value derived from energy meter readings
Epic Sounds GoPRo 37 47 Kitchens Audio labeled by external annotator
Appendix C Formal Definition of Dirichlet Distribution
Dirichlet distribution is often used to simulate data heterogeneity across different clients in FL. Dirichlet distribution is defined as follows: Given a K-dimensional vector $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},...,\alpha_{K})$ where each $\alpha_{i}>0$, a random vector $\boldsymbol{X}=(X_{1},X_{2},...,X_{K})$ follows a Dirichlet distribution, denoted as $\boldsymbol{X}\sim\text{Dir}(\boldsymbol{\alpha})$, if its probability density function (PDF) is given by:
$f(\boldsymbol{x};\boldsymbol{\alpha})=\frac{1}{\text{B}(\boldsymbol{\alpha})}\prod_{i=1}^{K}x_{i}^{\alpha_{i}-1}$
subject to the conditions $x_{i}\geq 0$ for all $i$ and $\sum_{i=1}^{K}x_{i}=1$. Here, $\text{B}(\boldsymbol{\alpha})$ is the multinomial beta function, defined as:
$\text{B}(\boldsymbol{\alpha})=\frac{\prod_{i=1}^{K}\Gamma(\alpha_{i})}{\Gamma\left(\sum_{i=1}^{K}\alpha_{i}\right)}$
where $\Gamma(\cdot)$ is the gamma function.
Appendix D Model Architectures
We evaluated multiple model architectures for each dataset and selected the one that performed the best. Table 11 lists the models examined for each dataset with the selected model marked in bold.
D.1 WISDM
For WISDM, we use a custom LSTM model that consists of an LSTM layer followed by a feed-forward neural network. The LSTM layer has an input dimension of 6 and a hidden dimension of 6. After the LSTM layer, the output is flattened and passed through a dropout layer with a rate of 0.2 for regularization. It then goes through a fully connected linear layer with an input size of $1,200$ (6 hidden units * 200 timesteps) and an output size of 128, followed by a ReLU activation function. Another dropout layer with a rate of 0.2 is applied before the final fully connected linear layer with an input size of 128 and an output size of 12.
D.2 UT-HAR
For UT-HAR, we use a ResNet-18 model with custom architecture designed for the Wi-Fi based Human Activity Recognition (HAR) task. The model consists of an initial convolutional layer that reshapes the input into a 3-channel tensor followed by the main ResNet architecture with 18 layers. This main architecture includes a series of convolutional blocks with residual connections, Group Normalization layers, ReLU activations, and max-pooling. Finally, there is an adaptive average pooling layer followed by a fully connected layer that outputs the class probabilities. The model utilizes 64 output channels in the initial layer and doubles the number of channels as it goes deeper. The last fully connected layer has 7 output units corresponding to the number of classes for the UT-HAR task.
D.3 Widar
For Widar, we also use a custom ResNet-18 model tailored for the Widar dataset. The model starts by reshaping the 22-channel input to 3 channels using two convolutional transpose layers, followed by a convolutional layer with 64 filters, Group Normalization, ReLU activation, and max-pooling. The core of the model consists of four layers of residual blocks (similar to the standard ResNet18) with 64, 128, 256, and 512 filters. Each basic block within these layers contains two convolutional layers, Group Normalization, and ReLU activations. Finally, an adaptive average pooling layer reduces spatial dimensions to $1\times 1$, followed by a fully connected layer to output class scores.
D.4 VisDrone
For VisDrone, we use the default YOLOv8n model from Ultralytics library. YOLOv8n is the smallest YOLOv8 model variant with the three scale parameters: depth, width, and the maximum number of channels set to 0.33, 0.25, and 1024 respectively.
D.5 CASAS
For CASAS, we use a BiLSTM neural network which is composed of an embedding layer, a bidirectional LSTM, and a fully connected layer. The embedding layer takes input sequences with dimensions equal to the input dimension and converts them to dense vectors of size 64. The bidirectional LSTM layer has an input size equal to 64, the same number of hidden units, and processes the embedded sequences in both forward and backward directions. The output of the LSTM layer is connected to a fully connected layer with an input size of 128 (to account for the bidirectional LSTM concatenation) and outputs the logits for 12 activities in the CASAS dataset.
D.6 AEP
For AEP, we use a custom multi-layer perceptron (MLP) neural network with an architecture comprising five hidden layers and an output layer. The input layer accepts 18 features and passes them through a linear transformation to the first hidden layer with 210 units. Each of the following hidden layers progressively scales the number of units by factors of 2 and 4 and then scales down. Specifically, the sizes of the hidden layers are 210, 420, 840, 420, and 210 units respectively. Each hidden layer uses a ReLU activation function followed by a dropout layer with a dropout rate of 0.3 for regularization. The output layer has a single unit, and the output of the network is obtained by passing the activations of the last hidden layer through a final linear transformation.
D.7 EPIC-SOUNDS
For EPIC-SOUNDS, we again use a custom ResNet-18 model which consists of a stack of convolutional layers followed by batch normalization and ReLU activation. The architecture begins with a $7\times 7$ convolutional layer with stride 2, followed by a max pooling layer. Then, it contains four blocks, each comprising a sequence of basic blocks with a residual connection; specifically, each block contains two basic blocks, with output channel sizes of 64, 128, 256, and 512 respectively. Each basic block comprises two sets of 3x3 convolutional layers, each followed by batch normalization and ReLU activation. The first convolutional layer in the basic block has a stride of 2 in the second, third, and fourth blocks. Finally, the model has an adaptive average pooling layer, which reduces the spatial dimensions to $1\times 1$, followed by a fully connected layer with an output size of 44 classes.
Table 11: List of models examined for each dataset. Selected model marked in bold.
{tblr}
width = colspec = Q[133]Q[158]Q[158]Q[120]Q[138]Q[142]Q[117], cells = c, hlines, WISDM & UT-HAR WIDAR VisDrone Casas AEP Epic Sounds
LSTM,
GRU
3-layer 1D CNN 3-layer Dense,
LeNet,
Resnet-18,
LSTM,
GRU,
BiLSTM,
3 layer 1D CNN,
2 layer Transformer 3-layer Dense,
LeNet,
Resnet-18,
LSTM,
GRU,
BiLSTM,
3 layer 1D CNN,
2 layer Transformer Yolo v3,
Yolo v5,
Yolo v8 LSTM,
BiLSTM,
3-layer LSTM,
3-layer BiLSTM 5-layer MLP,
SVM,
Linear Regressor ResNet18,
VGG-16
Appendix E Hyperparameters
Hyperparameters for Table 5. For WISDM-W, the learning rate for centralized training was 0.01 and we trained for 200 epochs with batch size 64. For FedAvg, in both low and high data heterogeneity scenarios, we used a client learning rate of 0.01 and trained for 400 communication rounds with batch size 32. For FedOPT, in both low and high data heterogeneity scenarios, we used a client learning rate of 0.01 and a server learning rate of 0.01. We also trained for 400 communication rounds. For WISDM-P, the learning rate for centralized training was 0.01 and we trained for 200 epochs with batch size 128. For FedAvg, in both low and high data heterogeneity scenarios, we used a client learning rate of 0.008 and trained for 400 communication rounds with batch size 32. For FedOPT, in both low and high data heterogeneity scenarios, we used a client learning rate of 0.01 and a server learning rate of 0.01. We also trained for 400 communication rounds. For UT-HAR and Widar, the learning rate for centralized training was 0.001 and the number of epochs was 500 and 200 for UT-HAR and Widar respectively with a batch size of 32. For both low and high data heterogeneity in both FedAvg and FedOPT, the client learning rate was 0.01 and the server learning rate for FedAvg and FedOPT was 1 and 0.01 respectively. The number of communication rounds was 1200 and 900 for UT-HAR and Widar respectively with a batch size of 32. For VisDrone, we used a cosine learning rate scheduler with $T_{0}=10,T_{mult}=2$ and trained for 200 epochs with a learning rate of 0.1 and batch size 12. For all the experiments on VisDrone, the client learning rate was also 0.1 and the batch size was 12. For FedOPT, the server learning rate was 0.1. For CASAS, the centralized learning rate was 0.1 with batch size 128. For the federated setting, the client learning rate was 0.005, and the batch size was 32. We trained for 400 rounds. For FedOPT, the server learning rate was 0.01. For AEP, the learning rate for centralized training was 0.001 and the batch size was 32 and it was trained for 1200 epochs. For federated experiments, the client learning rate was 0.01, and the batch size was 32. For FedOPT, the server learning rate was 0.1. For EPIC-SOUNDS, for centralized training, the learning rate was 0.1 with batch size 512. The number of epochs was 120. For federated settings, we used a client learning rate of 0.1 and batch size 32. For FedOPT, the server learning rate was 0.01.
Hyperparameters for Table 6. The setup for all the datasets with $10$% client sampling rate is the same as that of Table 5 under high data heterogeneity. For the $30$% client sampling rate, the hyperparameters were kept the same as that of the $10$% client sampling rate experiments, with the exception of CASAS, where the learning rate was set to 0.15.
Hyperparameters for Table 7. The hyperparameters were the same as that of Table 5 with $10$% sampling rate under high data heterogeneity scenario.
Hyperparameters for Table 8. The hyperparameters were same as that of Table 5 with $10$% client sampling rate under high data heterogeneity scenario.